OpenStudy (anonymous):

sam purchased 625 square feet of carpet for his bedroom which is shaped as a square what are the dimensions of his bedroom

OpenStudy (campbell_st):

just take the square root of 625 to get the dimensions... the room is square, all sides equal so the area is A = l^2 so 625 = l^2 so just take the square root

OpenStudy (anonymous):

wait I don't get it I am sorry would you explan it in an other way plz :)

OpenStudy (campbell_st):

ok... so how do you find the area of a square...?

OpenStudy (anonymous):

I am not that sure

OpenStudy (campbell_st):

ok.... well that makes it extremely difficult to answer this question have a look at this link about area. It may help http://www.mathopenref.com/area.html

OpenStudy (anonymous):

ooo ok ya I am kinda getting it amm I do know how to find the square root but I just don't know how to explain it